I need help please.
It is a velocity problem:
An aviator is timed going both directions over a 90 mile course. The time eastward, with the wind, is 24 minutes and westward, against the wind, is 30 minutes. Find the velocity of the wind in miles per hour.
Thank you in advance for your help.

Answer by scott8148(6628)

t = d / r

24/60 = 90 / (p + w) ___ 2p + 2w = 450

30/60 = 90 / (p - w) ___ p - w = 180 ___ p = 180 + w

substituting ___ 2(180 + w) + 2w = 450 ___ 4w + 360 = 450